A car is moving at a velocity of 10kmh-1 , on a straight road. The driver of the car throws a parcel with a velocity of 10√2 kmh-1, whne the car is passing by a man standing on the side of the road. If the parcel is to reach the man the direction of the throw makes the angle with the man?
 

Grade:11

2 Answers

Arun
4 years ago
Dear student
 
sin theta = 10/ 10sqrt(2) = 1/ sqrt(2)
 
hence theta = 45 degree
 
hence complete angle = 90 + 45 = 135 degree
 
Khimraj
4 years ago
Let the parcel be thrown at angle \Theta with the direction of car’s velocity
velocity of particle wrt ground
v = 10 i + 10{\sqrt{2}}(cos\Theta i + sin\Theta j)
if  the parcel is to reach the man 
10 + 10{\sqrt{2}}cos\Theta = 0
\Theta = 135